OpenStudy (anonymous):

6 = a/4 + 2 PLEASE HELP! THANKS! :) MEDAL

OpenStudy (ccswims):

You must cancel out 2 by subtracting So do 6-2

OpenStudy (anonymous):

4

OpenStudy (ccswims):

so now it's 4 = a/4 You need to get the variable by itself so do the inverse operation Do 4*4

OpenStudy (ccswims):

so the final answer is a = 16

OpenStudy (anonymous):

a = 1 because 4/1 = 4 and 4+2 =6

OpenStudy (ccswims):

Checking our work we see that 16/4 = 4 +2 = 6. So 16 is correct

OpenStudy (anonymous):

Thanks! @ccswims

OpenStudy (anonymous):

did i do something wrong?